# Dusit Niyato, , Bo Ai

### v1

@@ IV-B DG-OBE Combining Scheme Design
$\begin{cases}\mathbf{0},\,\,\mathrm{if}\,\,l\in\mathcal{P}_{k}\backslash\left\{k\right\}\,\,\\ p_{k}^{2}\tau_{p}\bar{\mathbf{g}}_{mkk}\tilde{\mathbf{r}}_{mkk}^{H}+p_{k}^{2}\tau_{p}\tilde{\mathbf{r}}_{mkk}\bar{\mathbf{g}}_{mkk}^{H},\mathrm{if}\,\,l=k\\ \end{cases}$
with $\tilde{\mathbf{r}}_{mlk}=\mathrm{vec}(\check{\mathbf{R}}_{ml}\mathbf{\Psi}_{mk}^{-1}\check{\mathbf{R}}_{mk})$ and $\bar{\mathbf{g}}_{mkk}=\mathrm{vec}(\bar{\mathbf{G}}_{mkk})$, and $\mathbf{\Lambda}_{mk}^{(4)}=\bar{\mathbf{R}}_{mk}^{T}\otimes\mathbf{I}_{N}$. Moreover, we have $\mathbf{\Lambda}_{kl}^{\left(3\right)}=p_{l}(\check{\mathbf{B}}_{lk}-\bar{\mathbf{B}}_{lk})$, where $\check{\mathbf{B}}_{lk}=\mathbf{b}_{lk}\mathbf{b}_{lk}^{H}$ and $\bar{\mathbf{B}}_{lk}=\mathrm{diag}(\bar{\mathbf{B}}_{1lk},\dots,\bar{\mathbf{B}}_{Mlk})\in\mathbb{C}^{MN^{2}\times MN^{2}}$ with $\mathbf{b}_{lk}=\left[\mathbf{b}_{1lk}^{T},\dots,\mathbf{b}_{Mlk}^{T}\right]^{T}$, $\bar{\mathbf{B}}_{mlk}=\mathbf{b}_{mlk}\mathbf{b}_{mlk}^{H}$, and $\mathbf{b}_{mlk}=\mathrm{vec}(\mathbf{B}_{mlk})$. The maximum value of SINR can also be computed in closed-form as $\overline{\widetilde{\mathrm{SINR}}}_{k}^{\mathrm{LSFD},*}=p_{k}\overline{\mathbf{h}}_{k}^{H}\overline{\mathbf{w}}_{k}^{\mathrm{d},*}$.
@@
As observed in Corollary 2 and Theorem 4, the optimal design of $\mathbf{W}_{mk}$ relies on global channel statistics information. However, given that channel statistics remain constant over a long period of time, it is feasible to implement the optimal design of $\mathbf{W}_{mk}$ locally at each AP. To facilitate this, the necessary channel statistics can be transmitted from other APs to each AP via fronthaul links for each realization of the AP/UE locations. Each AP can implement the DG-OBE combining scheme by plugging $\mathbf{W}_{mk}^{*}$ into (13). It is worth noting that the DG-OBE combining scheme allows for the distributed implementation at each AP by utilizing $\mathbf{W}_{mk}^{*}$ based on global channel statistics and local instantaneous channel estimates $\hat{\mathbf{g}}_{mk}$ at each AP. The results in Theorem 4 can reduce to so-called “GMR” combining as [25, Eq. (15)] when $\bar{\mathbf{G}}_{mkl}=\mathbf{0}$ and $N=1$.
Note that the optimal design of $\mathbf{W}_{mk}$ is implemented based on the performance analysis framework in (26). It is also available for the LSFD processing scheme discussed in Sec. III-B by plugging the derived DG-OBE schemes into results in Sec. III-B. We notice that, when applying the BE combining scheme in (13) at each AP, the final decoding data in the LSFD processing scheme can be denoted by $\check{x}_{k}=(\sum_{m=1}^{M}{a_{mk}^{*}\hat{\mathbf{g}}_{mk}^{H}\mathbf{W}_{mk}^{H}\mathbf{g}_{mk}})x_{k}+\sum_{l\neq k}^{K}{\sum_{m=1}^{M}{a_{mk}^{*}\hat{\mathbf{g}}_{mk}^{H}\mathbf{W}_{mk}^{H}\mathbf{g}_{ml}x_{l}}}+\mathbf{n}_{k}^{\prime}$. The LSFD coefficient $a_{mk}$ can be regarded as an additional scaling factor for $\mathbf{W}_{mk}$.
@@
For the conventional LSFD processing scheme with the arbitrary $\mathbf{W}_{mk}$, the LSFD coefficients can be optimized as in (16). However, when DG-OBE scheme in Corollary 2 is applied at each AP, the final decoding step at the CPU using the optimal LSFD coefficients becomes unnecessary since it can be regarded as that BE matrices with additional LSFD coefficients $\tilde{\mathbf{W}}_{mk}=a_{mk}\mathbf{W}_{mk}$ have been unifiedly optimized to maximize the achievable SE in (14). So there will be no further SE performance improvement from applying the optimal LSFD strategy if the DG-OBE combining scheme in Corollary 2 is applied. This observation will also be substantiated in the simulation results, like Fig. 4. In other words, the optimal LSFD scheme design in (16) can be considered to be a special case of DG-OBE scheme, where $\mathbf{W}_{mk}$ is fixed except for a scaling factor. Moreover, the DG-OBE scheme in Corollary 2 reduces to the optimal LSFD scheme in (16) under the case with single-antenna APs.
@@ IV-C DL-OBE Combining Scheme Design
Note that the DG-OBE combining scheme introduced in Sec. IV-B is designed based on the global channel statistics information. To realize the distributed design at each AP, the required statistics information needs to be delivered from the other APs via fronthaul links. In practice, the user scheduling could change rapidly so then one anyway needs to transfer statistics often. To relieve the often statistics transfer when designing the OBE scheme, in this subsection, we study the DL-OBE combining scheme design. As evident from its name, only the local channel statistics information at each AP is needed to locally design this combining scheme.
@@
$\frac{p_{k}|\mathbb{E}\{\mathbf{v}_{mk}^{H}\mathbf{g}_{mk}\}|^{2}}{\sum_{l=1}^{K}{p_{l}\mathbb{E}\{|\mathbf{v}_{mk}^{H}\mathbf{g}_{ml}|^{2}\}}-p_{k}|\mathbb{E}\{\mathbf{v}_{mk}^{H}\mathbf{g}_{mk}\}|^{2}+\sigma^{2}\mathbb{E}\{\|\mathbf{v}_{mk}\|^{2}\}},$
which can be easily proven based on the method of the centralized processing scheme. Based on (LABEL:SINR_Local), we can derive the value of $\mathbf{W}_{mk}$, that maximizes (LABEL:SINR_Local) as follows.
@@
The optimal $\mathbf{W}_{mk}$, maximizing the effective SINR for UE $k$ at AP $m$ as (LABEL:SINR_Local), can be denoted by $\mathbf{W}_{mk}^{*}=\mathrm{vec}^{-1}(\mathbf{w}_{mk}^{*})$ with the maximum value $\mathrm{SINR}_{mk}^{\mathrm{d},*}=p_{k}\mathrm{vec}(\mathbb{E}\{\mathbf{g}_{mk}\hat{\mathbf{g}}_{mk}^{H}\})^{H}\mathbf{w}_{mk}^{*}$, where $\mathbf{w}_{mk}^{*}=[\sum_{l=1}^{K}{p_{l}\mathbb{E}\{(\hat{\mathbf{g}}_{mk}\hat{\mathbf{g}}_{mk}^{H})^{T}\otimes(\mathbf{g}_{ml}\mathbf{g}_{ml}^{H})\}}-p_{k}\mathrm{vec}(\mathbb{E}\{\mathbf{g}_{mk}\hat{\mathbf{g}}_{mk}^{H}\})\mathrm{vec}(\mathbb{E}\{\mathbf{g}_{mk}\hat{\mathbf{g}}_{mk}^{H}\})^{H}+(\mathbb{E}\{\hat{\mathbf{g}}_{mk}\hat{\mathbf{g}}_{mk}^{H}\}^{T}\otimes\mathbf{I}_{N})]^{-1}\mathrm{vec}(\mathbb{E}\{\mathbf{g}_{mk}\hat{\mathbf{g}}_{mk}^{H}\})$.
@@
$\overline{\mathbf{w}}_{mk}^{*}=\mathbf{\Gamma}_{mk}^{-1}\bar{\mathbf{r}}_{mk},$ (30)
where $\mathbf{\Gamma}_{mk}=\sum\nolimits_{l=1}^{K}{p_{l}[(\bar{\mathbf{G}}_{mkk}^{T}\otimes\bar{\mathbf{G}}_{mll})+(\bar{\mathbf{G}}_{mkk}^{T}\otimes\check{\mathbf{R}}_{ml})}+(\hat{\mathbf{R}}_{mk}^{T}\otimes\bar{\mathbf{G}}_{mll})+(\hat{\mathbf{R}}_{mk}^{T}\otimes\check{\mathbf{R}}_{ml})]+\sum\nolimits_{l\in\mathcal{P}_{k}}{p_{k}p_{l}^{2}\tau_{p}^{2}\tilde{\mathbf{r}}_{mlk}\tilde{\mathbf{r}}_{mlk}^{H}}+p_{k}^{2}\tau_{p}\bar{\mathbf{g}}_{mkk}\tilde{\mathbf{r}}_{mkk}^{H}+p_{k}^{2}\tau_{p}\tilde{\mathbf{r}}_{mkk}\bar{\mathbf{g}}_{mkk}^{H}-p_{k}\bar{\mathbf{r}}_{mk}\bar{\mathbf{r}}_{mk}^{H}+\sigma^{2}(\bar{\mathbf{R}}_{mk}^{T}\otimes\mathbf{I}_{N})\in\mathbb{C}^{N^{2}\times N^{2}}$. The closed-form result of $\mathrm{SINR}_{mk}^{\mathrm{d},*}$ is $\overline{\mathrm{SINR}}_{mk}^{\mathrm{d},\mathrm{UatF},*}=p_{k}\bar{\mathbf{r}}_{mk}^{H}\overline{\mathbf{w}}_{mk}^{*}$.
@@
TABLE I: Computational complexity of considered schemes for each realization of AP/UE locations.
@@
$\frac{p_{k}|\mathbf{v}_{mk}^{H}\hat{\mathbf{g}}_{mk}|^{2}}{\sum\limits_{l\neq k}^{K}{p_{l}|\mathbf{v}_{mk}^{H}\hat{\mathbf{g}}_{ml}|^{2}}+\mathbf{v}_{mk}^{H}(\sum\limits_{l=1}^{K}{p_{l}\mathbf{C}_{ml}}+\sigma^{2}\mathbf{I}_{N})^{-1}\mathbf{v}_{mk}},$
is the instantaneous effective SINR for UE $k$ at AP $m$. When this standard bound is applied, the L-MMSE combining as in (12) cam maximize (LABEL:SINR_Distributed_Standard). However, the L-MMSE combining is designed based on the instantaneous channel estimates. When the UatF capacity bound in (LABEL:SINR_Local) is applied, the DL-OBE combining scheme introduced in this part can maximize (LABEL:SINR_Local).
@@
When the DL-OBE combining scheme is applied at each AP, the distributed processing scheme in Sec. III-B with the LSFD processing scheme, instead of the fully local decoding scheme as in (LABEL:SINR_Local), can be implemented to further enhance the SE performance for the networks. Thus, the DL-OBE combining scheme can be regarded as a heuristic one since, from the perspective of the decoding scheme in the whole network. And the fully local decoding scheme as in (LABEL:SINR_Local) is not applied for the decoding but only for the DL-OBE combining scheme design.
In Table I, we summarize the computational complexity of the considered schemes in this paper, where $I_{r}$ denotes the number of channel realizations and $\lfloor K/\tau_{p}\rfloor$ denotes the average number of UEs sharing the same pilot signal. As observed, the OBE-based combining scheme design embraces much lower computational complexity than that of MMSE-based combining scheme design, albeit at the cost of higher precomputation complexity. However, this higher precomputation complexity is deemed acceptable since it relies solely on channel statistics and needs to be performed only once for each realization of AP/UE locations.
@@
Compared to MMSE combining schemes, which rely on matrix inversion based on instantaneous channel estimates, OBE schemes are less sensitive to the quality of channel estimates due to their reliance on channel statistics-based OBE matrices. Therefore, OBE schemes are expected to remain robust in scenarios characterized by poor-quality channel estimates or severe pilot contamination. This observation will be further substantiated through numerical results, like Fig. 2, Fig. 3, and Fig. 5.

### v2

@@ IV-B DG-OBE Combining Scheme Design
$\begin{cases}\mathbf{0},\,\,\mathrm{if}\,\,l\in\mathcal{P}_{k}\backslash\left\{k\right\}\,\,\\ p_{k}^{2}\tau_{p}\overline{\mathbf{g}}_{mkk}\tilde{\mathbf{r}}_{mkk}^{H}+p_{k}^{2}\tau_{p}\tilde{\mathbf{r}}_{mkk}\overline{\mathbf{g}}_{mkk}^{H},\mathrm{if}\,\,l=k\\ \end{cases}$
with $\tilde{\mathbf{r}}_{mlk}=\mathrm{vec}(\check{\mathbf{R}}_{ml}\mathbf{\Psi}_{mk}^{-1}\check{\mathbf{R}}_{mk})$, $\overline{\mathbf{g}}_{mkk}=\mathrm{vec}(\overline{\mathbf{G}}_{mkk})$, and $\mathbf{\Lambda}_{mk}^{(4)}=\overline{\mathbf{R}}_{mk}^{T}\otimes\mathbf{I}_{N}$. Moreover, we have $\mathbf{\Lambda}_{kl}^{\left(3\right)}=p_{l}(\check{\mathbf{B}}_{lk}-\overline{\mathbf{B}}_{lk})$, where $\check{\mathbf{B}}_{lk}=\mathbf{b}_{lk}\mathbf{b}_{lk}^{H}$ and $\overline{\mathbf{B}}_{lk}=\mathrm{diag}(\overline{\mathbf{B}}_{1lk},\dots,\overline{\mathbf{B}}_{Mlk})\in\mathbb{C}^{MN^{2}\times MN^{2}}$ with $\mathbf{b}_{lk}=\left[\mathbf{b}_{1lk}^{T},\dots,\mathbf{b}_{Mlk}^{T}\right]^{T}$, $\overline{\mathbf{B}}_{mlk}=\mathbf{b}_{mlk}\mathbf{b}_{mlk}^{H}$, and $\mathbf{b}_{mlk}=\mathrm{vec}(\mathbf{B}_{mlk})$. The maximum value of SINR can also be computed in closed-form as $\overline{\widetilde{\mathrm{SINR}}}_{k}^{\mathrm{LSFD},*}=p_{k}\overline{\mathbf{h}}_{k}^{H}\overline{\mathbf{w}}_{k}^{\mathrm{d},*}$.
@@
As observed in Corollary 2 and Theorem 4, the optimal design of $\mathbf{W}_{mk}$ relies on global channel statistics information. However, given that channel statistics remain constant over a long period of time, it is feasible to implement the optimal design of $\mathbf{W}_{mk}$ locally at each AP. To facilitate this, the necessary channel statistics can be transmitted from other APs to each AP via fronthaul links for each realization of the AP/UE locations. Each AP can implement the DG-OBE combining scheme by plugging $\mathbf{W}_{mk}^{*}$ into (13). It is worth noting that the DG-OBE combining scheme allows for the distributed implementation at each AP by utilizing $\mathbf{W}_{mk}^{*}$ based on global channel statistics and local instantaneous channel estimates $\widehat{\mathbf{g}}_{mk}$ at each AP. The results in Theorem 6 can reduce to so-called “GMR” combining as [28, Eq. (15)] when $\overline{\mathbf{G}}_{mkl}=\mathbf{0}$ and $N=1$.
Note that the optimal design of $\mathbf{W}_{mk}$ is implemented based on the performance analysis framework in (26). It is also available for the LSFD processing scheme discussed in Sec. III-B by plugging the derived DG-OBE schemes into results in Sec. III-B. We notice that, when applying the BE combining scheme in (13) at each AP, the final decoding data in the LSFD processing scheme can be denoted by $\check{x}_{k}=(\sum_{m=1}^{M}{a_{mk}^{*}\widehat{\mathbf{g}}_{mk}^{H}\mathbf{W}_{mk}^{H}\mathbf{g}_{mk}})x_{k}+\sum_{l\neq k}^{K}{\sum_{m=1}^{M}{a_{mk}^{*}\widehat{\mathbf{g}}_{mk}^{H}\mathbf{W}_{mk}^{H}\mathbf{g}_{ml}x_{l}}}+\mathbf{n}_{k}^{\prime}$. The LSFD coefficient $a_{mk}$ can be regarded as an additional scaling factor for $\mathbf{W}_{mk}$.
@@
For the conventional LSFD processing scheme with the arbitrary $\mathbf{W}_{mk}$, the LSFD coefficients can be optimized as in (16). However, when DG-OBE scheme in Corollary 2 is applied at each AP, the final decoding step at the CPU using the optimal LSFD coefficients becomes unnecessary since it can be regarded as that BE matrices with additional LSFD coefficients $\tilde{\mathbf{W}}_{mk}=a_{mk}\mathbf{W}_{mk}$ have been unifiedly optimized to maximize the achievable SE in (14). So there will be no further SE performance improvement from applying the optimal LSFD strategy if the DG-OBE combining scheme in Corollary 2 is applied. This observation will also be substantiated in the simulation results, like Fig. 5. In other words, the optimal LSFD scheme design in (16) can be considered to be a special case of DG-OBE scheme, where $\mathbf{W}_{mk}$ is fixed except for a scaling factor. Moreover, the DG-OBE scheme in Corollary 2 reduces to the optimal LSFD scheme in (16) with the L-MR combining under the case with single-antenna APs.
Remark 8.
Note that the performance analysis framework for the distributed processing scheme in (14) holds for arbitrary local combining vectors $\mathbf{v}_{mk}$. The distributed OBE scheme $\mathbf{v}_{mk}=\mathbf{W}_{mk}\widehat{\mathbf{g}}_{mk}$ introduced in this part can be equivalently viewed as the combination of the L-MR combining scheme $\overline{\mathbf{v}}_{mk}=\widehat{\mathbf{g}}_{mk}$ and the weighting matrix $\mathbf{W}_{mk}$. Furthermore, a more generalized decoding scheme can be extended from the distributed OBE scheme as $\mathbf{v}_{mk}=\mathbf{W}_{mk}\overline{\mathbf{v}}_{mk}$, where $\overline{\mathbf{v}}_{mk}$ denotes an arbitrary local combining vector. Following a similar approach as in Corollary 2, the weighting matrix $\mathbf{W}_{mk}$ can be optimized to maximize the achievable SE performance by replacing $\widehat{\mathbf{g}}_{mk}$ in (LABEL:DG_OBE_W_Monte) by $\overline{\mathbf{v}}_{mk}$, which can be called as the OBE weighting strategy. Different from Corollary 2, the optimal OBE weighting strategy holds for an arbitrary local combining vector $\overline{\mathbf{v}}_{mk}$. Moreover, different from the local channel estimates-based combining scheme, the channel state information from other APs can also be involved in the design of $\overline{\mathbf{v}}_{mk}$. One possible scheme is the team MMSE combining scheme, where each AP can involve the channel estimates from other APs in a sequential manner to design the local combining scheme as in [33]. By constructing these local combining schemes with the aid of the channel state information from other APs or neighboring APs, we can derive the corresponding OBE weighting strategy. This combination is anticipated to further enhance the SE performance compared to the scheme with the OBE weighting scheme and local combining scheme with only local channel state information at each AP.
@@ IV-C DL-OBE Combining Scheme Design
Note that the DG-OBE combining matrices introduced in Sec. IV-B is designed based on the global channel statistics information. To realize the distributed design at each AP, the required statistics information needs to be delivered from the other APs via fronthaul links. In practice, the user scheduling could change rapidly so then one anyway needs to transfer statistics often. To relieve the often statistics transfer when designing the OBE scheme, in this subsection, we study the DL-OBE combining scheme design. As evident from its name, only the local channel statistics information at each AP is needed to locally design the OBE matrices in this combining scheme.
@@
$\frac{p_{k}|\mathbb{E}\{\mathbf{v}_{mk}^{H}\mathbf{g}_{mk}\}|^{2}}{\sum\limits_{l=1}^{K}{p_{l}\mathbb{E}\{|\mathbf{v}_{mk}^{H}\mathbf{g}_{ml}|^{2}\}}-p_{k}|\mathbb{E}\{\mathbf{v}_{mk}^{H}\mathbf{g}_{mk}\}|^{2}+\sigma^{2}\mathbb{E}\{\|\mathbf{v}_{mk}\|^{2}\}},$
which can be easily proven based on the method of the centralized processing scheme. Based on (LABEL:SINR_Local), we can derive the optimal $\mathbf{W}_{mk}$, that maximizes (LABEL:SINR_Local) as follows.
@@
The optimal $\mathbf{W}_{mk}$, maximizing the effective SINR for UE $k$ at AP $m$ as (LABEL:SINR_Local), can be denoted by $\mathbf{W}_{mk}^{*}=\mathrm{vec}^{-1}(\mathbf{w}_{mk}^{*})$ with the maximum value $\mathrm{SINR}_{mk}^{\mathrm{d},*}=p_{k}\mathrm{vec}(\mathbb{E}\{\mathbf{g}_{mk}\widehat{\mathbf{g}}_{mk}^{H}\})^{H}\mathbf{w}_{mk}^{*}$, where $\mathbf{w}_{mk}^{*}=[\sum_{l=1}^{K}{p_{l}\mathbb{E}\{(\widehat{\mathbf{g}}_{mk}\widehat{\mathbf{g}}_{mk}^{H})^{T}\otimes(\mathbf{g}_{ml}\mathbf{g}_{ml}^{H})\}}-p_{k}\mathrm{vec}(\mathbb{E}\{\mathbf{g}_{mk}\widehat{\mathbf{g}}_{mk}^{H}\})\mathrm{vec}(\mathbb{E}\{\mathbf{g}_{mk}\widehat{\mathbf{g}}_{mk}^{H}\})^{H}+(\mathbb{E}\{\widehat{\mathbf{g}}_{mk}\widehat{\mathbf{g}}_{mk}^{H}\}^{T}\otimes\mathbf{I}_{N})]^{-1}\mathrm{vec}(\mathbb{E}\{\mathbf{g}_{mk}\widehat{\mathbf{g}}_{mk}^{H}\})$.
@@
$\overline{\mathbf{w}}_{mk}^{*}=\mathbf{\Gamma}_{mk}^{-1}\overline{\mathbf{r}}_{mk},$ (30)
where $\mathbf{\Gamma}_{mk}=\sum\nolimits_{l=1}^{K}{p_{l}[(\overline{\mathbf{G}}_{mkk}^{T}\otimes\overline{\mathbf{G}}_{mll})+(\overline{\mathbf{G}}_{mkk}^{T}\otimes\check{\mathbf{R}}_{ml})}+(\widehat{\mathbf{R}}_{mk}^{T}\otimes\overline{\mathbf{G}}_{mll})+(\widehat{\mathbf{R}}_{mk}^{T}\otimes\check{\mathbf{R}}_{ml})]+\sum\nolimits_{l\in\mathcal{P}_{k}}{p_{k}p_{l}^{2}\tau_{p}^{2}\tilde{\mathbf{r}}_{mlk}\tilde{\mathbf{r}}_{mlk}^{H}}+p_{k}^{2}\tau_{p}\overline{\mathbf{g}}_{mkk}\tilde{\mathbf{r}}_{mkk}^{H}+p_{k}^{2}\tau_{p}\tilde{\mathbf{r}}_{mkk}\overline{\mathbf{g}}_{mkk}^{H}-p_{k}\overline{\mathbf{r}}_{mk}\overline{\mathbf{r}}_{mk}^{H}+\sigma^{2}(\overline{\mathbf{R}}_{mk}^{T}\otimes\mathbf{I}_{N})\in\mathbb{C}^{N^{2}\times N^{2}}$. The closed-form result of $\mathrm{SINR}_{mk}^{\mathrm{d},*}$ is $\overline{\mathrm{SINR}}_{mk}^{\mathrm{d},\mathrm{UatF},*}=p_{k}\overline{\mathbf{r}}_{mk}^{H}\overline{\mathbf{w}}_{mk}^{*}$.
@@
TABLE I: Computational complexity of considered schemes for each realization of the AP/UE locations.
@@
$\frac{p_{k}|\mathbf{v}_{mk}^{H}\widehat{\mathbf{g}}_{mk}|^{2}}{\sum\limits_{l\neq k}^{K}{p_{l}|\mathbf{v}_{mk}^{H}\widehat{\mathbf{g}}_{ml}|^{2}}+\mathbf{v}_{mk}^{H}(\sum\limits_{l=1}^{K}{p_{l}\mathbf{C}_{ml}}+\sigma^{2}\mathbf{I}_{N})^{-1}\mathbf{v}_{mk}},$
is the instantaneous effective SINR for UE $k$ at AP $m$. When this standard bound is applied, the L-MMSE combining in (12) can maximize (LABEL:SINR_Distributed_Standard). However, the L-MMSE combining is designed based on the instantaneous channel estimates. When the UatF capacity bound in (LABEL:SINR_Local) is applied, the DL-OBE combining scheme introduced in this part can maximize (LABEL:SINR_Local).
@@
When the DL-OBE combining scheme is applied at each AP, the distributed processing scheme in Sec. III-B with the LSFD processing scheme, instead of the fully local decoding scheme in (LABEL:SINR_Local), can be implemented to further enhance the SE performance for the networks. Thus, the DL-OBE combining scheme can be regarded as a heuristic one since, from the perspective of the decoding scheme in the whole network. And the fully local decoding scheme in (LABEL:SINR_Local) is not applied for the decoding but only for the DL-OBE combining scheme design.
In Table I, we summarize the computational complexity of the considered schemes in this paper, where $I_{r}$ denotes the number of channel realizations for each realizations of the AP/UE locations and $\lfloor K/\tau_{p}\rfloor$ denotes the average number of UEs sharing the same pilot signal. As observed, the OBE-based combining scheme design embraces much lower computational complexity than that of MMSE-based combining scheme design, albeit at the cost of higher precomputation complexity. This higher precomputation complexity is deemed acceptable since it relies solely on channel statistics and needs to be performed only once for each realization of the AP/UE locations.
However, in large-scale or highly dynamic networks, this precomputation complexity will become challenging. More specifically, for the large-scale or dense networks with large values of $\{M,K,N\}$, the precomputation complexity becomes quite large even if the computation occurs only once per major change in the AP/UE configuration. For highly dynamic networks with high user mobility or rapid changes in the AP/UE configurations, the channel statistics vary frequently, necessitating more frequent recalculations of the required precomputed matrices than those of the low mobility scenarios. These two important issues pose noticeable challenges for the practical implementation of the OBE schemes, especially the C-OBE scheme with the highest precomputation complexity. To facilitate the practical implementation of the proposed OBE schemes, here are three promising enablers:
(1) The distributed OBE schemes are advocated in large-scale or highly dynamic networks: As will be comprehensively discussed in the simulation part, such as Fig. 10, the distributed OBE schemes, especially the DL-OBE scheme, show competitive advantage in the highly dynamic scenario with much smaller precomputation complexity than the C-OBE scheme. Thus, to balance the achievable performance and precomputation complexity, the distributed OBE schemes, especially the DL-OBE scheme, are highly advocated in the large-scale or highly dynamic networks.
(2) Low-complexity algorithms can be applied to enhance the feasibility of OBE schemes in practice: Another promising solution is to apply some low-complexity algorithms to significantly reduce the precomputation complexity, such as the low-complexity matrix inversion algorithms. Taking the C-OBE scheme as an example, in (24), for each UE, the major computational complexity of the C-OBE scheme comes from the $M^{2}N^{2}\times M^{2}N^{2}$ matrix inversion, which is $\mathcal{O}(M^{6}N^{6})$. By applying low-complexity matrix inversion algorithms, such as the symmetric successive over relaxation (SSOR) algorithm [34], which can provide an approximate calculation of the matrix inversion, the computational complexity of this matrix inversion can be efficiently degraded to $\mathcal{O}(M^{4}N^{4}N_{Iter})$ with $N_{Iter}$ being the number of iterations of the SSOR algorithm, which is usually set to a small value. Besides, other low-complexity algorithms are also advocated to further degrade the precomputation complexity, which is regarded as an important future direction.
(3) Machine learning can be utilized to empower the OBE design in highly dynamic networks: To facilitate the OBE design in highly dynamic networks, where the statistics frequently vary due to the mobility and AP/UE configuration, the machine learning methodology can be utilized to help learn how the statistical parameters change in practice. Besides, another promising methodology is to learn how to estimate the statistics from received data signals, where the channel map can be applied to identify statistical parameters with data collected from previous users located on the same location [35]. Thus, the intelligent OBE design can be efficiently empowered by the machine learning methodology based on the learned statistical parameters. Moreover, by learning the vectorization and vectorization reversion approaches for the optimization of BE matrices, the machine learning is also anticipated to directly construct the OBE matrices based on the statistics but avoid the computation of OBE vectors with high computational complexity.
@@
All the results are derived for Rician fading channels, but they also hold in the special case of Rayleigh fading channels. More specifically, by letting $\overline{\mathbf{g}}_{mk}=\mathbf{0}$, all results reduce to those with Rayleigh fading. Notably, Fig. 8 verifies the results for the Rayleigh fading channel.
Remark 11.
Compared to MMSE combining schemes, which rely on matrix inversion based on instantaneous channel estimates, OBE schemes are less sensitive to the quality of channel estimates due to their reliance on channel statistics-based OBE matrices. Therefore, OBE schemes are expected to remain robust in scenarios characterized by poor-quality channel estimates or severe pilot contamination. This observation will be further substantiated through numerical results, like Fig. 3, Fig. 4, Fig. 6, and Fig. 9.